Give the examples of compact and noncompact curves in the plane R^2.


1
Expert's answer
2021-07-05T16:56:56-0400

To give the examples we can use the following fact: the compact subsets of "\mathbb{R}^n" are exactly the bounded closed sets.

Therefore the examples of compact curves would be:

  • finite segments "[a;b]" for two points "a,b\in\mathbb{R}^2",
  • circles "(x-a)^2+(y-b)^2=r^2" ,
  • finite segments of curves, i.e., "\mathcal{C}\cap B_{closed}(0,r)" for any closed curve "\mathcal{C}" and any "r>0",
  • an image of "[0;1]\subseteq \mathbb{R}" by any continuous function "f:[0;1]\to \mathbb{R}^2".

And the examples of noncompact curves would respectively be:

  • straight lines (such as "y=kx+b"),
  • parabolas "y=ax^2+bx+c" with "a\neq 0" , hyperbolas "(y-a)^2-(x-b)^2=c",
  • any other unbounded curves (sinusoids, curves "y=P(x)" with a non-constant polynomial "P" etc.)
  • not closed curves (circle minus a point, an open-ended ray, semi-open interval, open-ended branch of a parabola, etc.)